OpenStudy (anonymous):

Solve the given equation and indicate the number of values of θ sach that 0≤θ<2π,that satisfy the iquation ; 2sin^2θ-sinθ-1=0

OpenStudy (sgadi):

Take \[\sin\theta=x\] given equation \[2x^2-x-1=0\] solving above equation, we get x=-0.5 or x=1 or we can say \[\sin\theta = -0.5\] or \[\sin\theta = 1\] so we can say the values of theta for which this is true is \[\theta = \frac{7\pi}{6},\frac{11\pi}{6},\frac{3\pi}{2}\]